# Strain-Enhanced Altermagnetism in Ca3Ru2O7

### v1

@@
Abstract
Perovskite compounds exhibit enhanced octahedral distortions coupled with strong electronic correlations, providing a promising platform to explore and tune altermagnetic (AM) order. In this work, we investigate AM phases in Ca3Ru2O7, a well-known perovskite that hosts antiferromagnetism coupled to structural degrees of freedom. We demonstrate that Ca3Ru2O7 is a Kramers antiferromagnet in its ground state. However, a Néel-type magnetic configuration reveals a P-2 d-wave AM, hosting orbital-selective altermagnetism analogous to Ca2RuO4. We further explore the effects of biaxial strain on the stability between the antiferromagnetic ground state and the AM phase. Our results suggest that the AM phase becomes more stable than the AFM phase below -2% compressive strain. Additionally, strain can either enhance or suppress AM bands, with enhancements reaching up to 9% under tensile strain. To quantify the AM band splitting, we introduce the ”Altermagnetic Merit Figure” and analyze the role of electronic localization, delocalization, and octahedral distortions in AM behavior and magnetic stability changes under strain.
1 Introduction
[…]
Due to the breaking of time-reversal symmetry, the spin-orbit interaction can induce the so-called weak ferromagnetism [6] in altermagnet via a canting of the spins or other mechanisms depending on the crystal symmetries of the material [7, 8, 9, 10, 11, 12].
Given these findings, it is natural to explore whether strain could also play a key role in tuning AM order in other correlated materials. In this regard, Ruddlesden-Popper perovskites, which host diverse quantum states, could represent an interesting class of systems for investigating strain-induced AM behavior. Since many of their electronic phases are closely linked to structural distortions, these perovskites offer a promising platform to study the emergence of AM phases in relation to their inherent structural and electronic degrees of freedom [13, 14, 15, 16].
Ruddlesden-Popper compounds of the form (Ca,Sr)n+1RunO3n+1 are well-known correlated systems whose intriguing electronic properties are strongly influenced by their symmetries and octahedral distortions. These distortions, including rotations, Jahn-Teller effects, and polar distortions, can break symmetries that transform opposite spin sublattices, potentially enabling the emergence of AM states [14]. Over the past decade, these materials have attracted significant attention due to their exotic electronic phases, including Mott insulating behavior [17], metamagnetism [18], and unconventional superconductivity [19]. Revisiting their magnetic phases in the context of AM order could provide new insights into phase transitions between ferromagnetism, antiferromagnetism, and AM, offering a framework to understand the interplay between symmetry, octahedral distortions, and magnetism in correlated systems.
Among the (Ca,Sr)n+1RunO3n+1 series, we focus on Ca3Ru2O7 (CRO) to explore AM states. CRO is a fascinating compound that undergoes magnetic AFM transitions with temperature [20, 21, 22]. Over the past decade, experiments have shown that its electronic phases can be manipulated by magnetic fields [23], pressure [24], and, more recently, strain [25]. This raises the question of whether CRO exhibits altermagnetism or under what conditions its AFM properties can be tuned toward an AM state. Moreover, very few altermagnetic materials lack inversion symmetry[26, 27]. Tuning altermagnetism in Ca3Ru2O7 would introduce another noncentrosymmetric AM material, enabling the study of the interplay between non-relativistic and relativistic spin-splitting.
In this work, we explore various magnetic configurations in the CRO system to identify those that exhibit AM order. We begin by analyzing the electronic properties and then discuss the symmetry characteristics that give rise to AM phases in CRO. Furthermore, we examine the system under strain to determine the conditions under which the AM phase can be stabilized. Our results demonstrate that the altermagnetic character can be enhanced or suppressed under biaxial tensile and compressive strain. We discuss the role of octahedral distortions and the localization/delocalization effects induced by the strain in the changes to altermagnetic band splitting. Additionally, we introduce the altermagnetic quantity (AMQ) as the “figure of merit” to quantify the degree of enhancement or suppression of the band splitting with respect to the non-strained system.
2 Computational details
We perform a theoretical analysis using the density functional theory with and without spin-orbit coupling (SOC). We use the plane-wave pseudopotential method implemented in the Vienna ab-initio simulation package (VASP)[28] within the generalized gradient approximations (GGA), and it gives structural properties closer to the experimental values. The electronic valence considered are: Ru: 54 and O: 22, and Ca: 334. We use a plane-wave energy cutoff of 650 eV and set a Regular Monkhorst-Pack grid of 775 to perform the atomic relaxation and 11117 to perform the self-consistent calculation. We use a fine k-grid of 14147 within the tetrahedron method for the density of states. We perform the structural optimization of the unit cell until a force convergence threshold of at least 10-3 eV/Å per atom.
To account for the electronic correlation effects in the -orbitals of Ru atoms, we apply a Hubbard on-site Coulomb parameter using the Dudarev approximation [29] with U = 1.0 eV, a sufficiently accurate value for computing the electronic and structural properties of CRO, as reported in our previous study [30].
3 Results
3.1 Exploring AM states
@@
CRO exhibits a non-centrosymmetric crystal structure with the space group Bb21m [21]. The associated point group is mm2, which includes a twofold rotation axis along the -axis (C2z), two perpendicular mirror planes parallel to the and planes, and a helicoidal translation along (E, C2z, , , 21). This point group has been associated with the exhibition of altermagnetic (AM) phases [31]. The magnetic moment arises predominantly from the Ru 4 electrons, in which, at low temperatures, the Ru atoms order ferromagnetically within the bilayers while aligning antiferromagnetically between them, forming an A-type antiferromagnetic structure where the spins are oriented along the -axis (see Fig. 1). Given this A-type antiferromagnetic order, the primitive cell of CRO contains spin-up sites, with a translational symmetry operation that maps spin-up into spin-down sites (along the -axis). Consequently, the ground state is not altermagnetic (see Fig. 1) [32].
@@
To explore possible AM states and their stability, we analyzed CRO under different magnetic configurations. Specifically, we examined magnetic orderings commonly observed in AB3 perovskites [34], labeled as configurations A, B, C, and D (see Fig.2). Table 1 shows the energy differences relative to the ground-state configuration A, as well as some relevant structural, electronic, and magnetic properties for our study. Our results indicate that AM states appear in configurations B and C, where configuration B has an antiferromagnetic spin arrangement along the - and -axes, while configuration C shows this along the -axis. Although these magnetic states have not been experimentally observed in CRO, similar phases have been reported in doped Ca3Ru2O7 systems, such as Ti-doped Ca3(Ru1-xTix)2O7 (configuration B) [35], and in the sister compound Sr3Ru2O7 (configuration D) [36].
@@
Notably, variations in magnetic spin arrangements lead to significant structural and electronic changes (further structural details are given in SM table S1). Specifically, FM and AFM in-plane arrangements result in metallic (A, C) and a narrow insular state (B, D), respectively. Next, we will focus on the AM phases of CRO, primarily in Configuration B.
E | E-State | M-Phase | () | |
A | 0 | Metallic | AFM | 1.44 |
B | 34.90 | N-Insulator | AM | 1.35 |
C | 36.46 | Metallic | AM | 1.38/1.45 |
D | 39.22 | N-Insulator | AFM | 1.35 |
@@
In addition to the electronic differences, configuration B exhibits a lower magnetic moment than configuration A (details on Table 1).
In configuration B, the magnetic moments adopt a Néel-type ordering, where spin-up and spin-down orientations alternate within the -plane and along the -axis, forming spin-singlet wave functions.
@@
This configuration promotes spatial delocalization, which is characteristic of bonding-like states. Since bonding states enhance wave function overlap, that is, increase the spatial proximity of electrons, configuration B is more sensitive to the Hubbard- term.
@@
As a result, a band gap opens at relatively low U values, specifically . In contrast, configuration A remains metallic under the same conditions. Further analysis is provided in the Supplemental Material (Section SII).
@@
We now focus on the electronic band structure of both systems. Figure 3 shows the total and partial density of states for configurations A and B. Configuration A exhibits metallic behavior, while configuration B displays a narrow gap at the Fermi level. The Ru ions, located in an octahedral environment formed by oxygen atoms, are subjected to a crystal field that splits the Ru 4 levels, raising the eg states and lowering the t2g manifold energies. In configuration B, the orbitals are nearly filled, leading to a small gap in the half-filled and bands. Notably, the orbitals are the most affected, as the spin rearrangement occurs within the -plane. This behavior is similar to that observed in Ca2RuO4
[40].
@@
Figure 4 shows the non-relativistic band structure for the A and B magnetic configurations. The A configuration (ground state) exhibits metallic behavior, consistent with previously reported results that neglect spin-orbit coupling (SOC) effects [41, 30]. In contrast, configuration B displays a semimetallic character and a -dependent non-relativistic spin splitting, a hallmark of altermagnetic states (see Fig.4(b)). The shaded regions on the band structure indicate regions in -space where the altermagnetic spin splitting is maximal. The detailed BZ for these paths is shown on the right Fig. 4.
@@
The presence of the non-relativistic spin-splitting is independent of the value of ky. According to the notation on the spin-momentum locking, this is P-2 d-wave altermagnets (where the d-wave is in the k-space) [32], since it is independent on ky this is a P-2 dxz-wave. These paths, which extend beyond the plane, have not been explored in previous studies, where ARPES experiments primarily focused on the plane [41]. In the next section, we discuss the preference for a specific direction in the AM character.
@@
The bands located between eV exhibit significant non-relativistic spin splitting, corresponding to the - orbitals. In contrast, the bands between eV, dominated by the orbital character, showing negligible spin splitting (see, Fig.LABEL:fig:S2). Therefore, configuration B exhibits orbital-selective altermagnetism with the altermagnetism present in the - bands but not in the bands, this is similar to that observed in Ca2RuO4 [40]. When the or coordinate is inverted in reciprocal space, the non-relativistic spin splitting changes sign, while inversion of leaves it unchanged. Notably, this altermagnetic order matches that found in Ca2RuO4, despite its crystallization in a different space group [40]. The orbital-selective magnetism depends on the magnetic space group and orbital character. In conf.B, the nearly filled dxy orbitals do not contribute to magnetism, leading to orbital-selective AM. Consequently, conf. C displays different magnetic states in which all the t2g orbitals contribute to the AM character; therefore, it does not exhibit orbital selective AM, see more details in Fig.LABEL:fig:S2.
CRO, due to its non-centrosymmetric structure, is expected to exhibit pronounced SOC effects, as demonstrated in previous studies on configuration A (the experimental magnetic phase), where the spins are oriented along the -axis [41, 30]. In configuration B, the anisotropy also lies primarily along the -axis, with a minor component along the -axis, specifically . Figure LABEL:fig:S3 in the SM shows the band structure with SOC, by spin-component contributions and orbitals. Since the magnetic moment aligns along the -axis, a strong Kramers spin degeneracy is evident along this direction. However, spin polarization components are also observed along the -axis. The main contributors to the AM character are the dxy orbitals (from -0.7 to -0.5 eV) and the dyz orbitals (from -1.5 to -1.0 eV) (see SM, Fig. LABEL:fig:S2). The slight enhancement of the dxy contribution arises from the band splitting in the band structure induced by SOC, on top of the non-relativistic spin splitting. This induces both splitting and mixing of the spin character of the bands.
3.2 Symmetry analysis
@@
In summary, the operation z becomes necessary (Fig. 5(f)). These combined nonsymmorphic operations result in the -dependent splitting states [32]. Due to the symmetry operations that connect the spins, which are determined by rotations and translations along the and directions while leaving the direction unchanged, the AM character is found along these axes, as shown in Fig. 4. In the case of the C configuration, the symmetry operation ( +, + , - + ) becomes necessary to achieve AM order.
3.3 Tuning AFM to AM phases
@@
Next, we focus on investigating the effects of compressive (-) and tensile (+) strain on the AM features of CRO. Fig. 7 shows the band structure of configuration B under strains of = -2%, 0, and 2% (biaxial). It can be observed that compressive strain suppresses the AM features among the energy range [-2:-1] eV (compare Fig. 7 (a) with Fig.7(b)). Moreover, the tensile strain enhances the splitting being shifted to higher energies.
@@
Note that the integration to obtain the AMQ can be performed either over the entire Brillouin zone or along a specific -path. We have verified that both methods yield similar solutions up to a constant multiplicative factor, which cancels out when percentage changes are considered.
We sum up all the occupied states rather than comparing only a few selected bands. This approach proves more reliable because the most strongly split bands can vary across different systems and even within the same material under strain, depending on details such as band character and crystal symmetry. Consequently, integrating over the entire set of occupied bands furnishes a more robust assessment of the altermagnetic spin-splitting.
@@
in our code, Simpson’s rule is used for numerical integration. When points are expressed in inverse angstroms (), the resulting AMQ is in units of . Furthermore, this straightforward procedure for quantifying the altermagnetic character may fail to correctly identify the spin-splitting in cases where bands cross or overlap. A potentially improved approach would incorporate a disentanglement method based on band character analysis, for instance, using Wannier functions or explicitly tracking the orbital character of each band. Such an approach would enable a more rigorous and accurate splitting assignment; however, its detailed implementation is beyond the scope of the present study and remains a subject for future work.
@@
Fig. S9 illustrates the color-map along the entire -path (see the BZ in Fig.4) and by occupied bands (number labels at the -axis) corresponding to each strain value. The splitting by bands clearly varies as a function of the applied strain, with some bands becoming more or less altermagnetic. Under tensile strain (up to 4% of the maximum value explored), increases significantly compared to the unstrained system. By integrating using Eq. 2, we obtain AQM values of 0.0010, 0.0012, and 0.0013 for strains of -2%, 0%, and 2%, respectively. Fig.6(c) shows the percentual variation of AQM across different strain values given by Eq. 3. Under biaxial strain, the band splitting increases by approximately under tensile strain (at 2%) and decreases by about under compressive strain (at -2%).
The non-relativistic spin-splitting in Ca3Ru2O7 arises from the interplay between hopping parameters and the on-site energy difference between majority and minority -electrons. Strain introduces competing effects that influence these parameters: compression along one axis promotes electronic delocalization, weakening magnetism and reducing the AM band splitting, whereas tensile strain favors localization, enhancing magnetic exchange and increasing the AM band splitting.
From the energy–strain relationship shown in Fig. 6(a), we observe that configuration B (the AM phase) becomes energetically more favorable than configuration A (the AFM phase) under biaxial compressive strain (). Compressive strain decreases the in-plane Ru-O bond lengths, forcing an expansion along the -axis. This expansion modifies the balance of structural distortions so that configuration B is stabilized relative to configuration A. At the electronic level, the in-plane compression broadens the bands, weakening the magnetic exchange interactions and consequently reducing the AM band splitting. Conversely, under tensile strain, the lattice is stretched in the -plane while slightly contracting along , effectively narrowing the bandwidth and reinforcing the altermagnetic character (see, Fig.S8 SM). This effect is reflected in the rising AMQ values in Fig. 6(c) RS line (relaxed structures), demonstrating how strain can selectively enhance or suppress altermagnetic character.

### v2

# Hybrid d/p-wave altermagnetism in Ca3Ru2O7 and strain-controlled spin splitting

@@
Hybrid d/p-wave altermagnetism in Ca3Ru2O7 and strain-controlled spin splitting
2Dresden University of Technology, Institute for Solid State and Materials Physics, 01062 Dresden, Germany
3International Research Centre Magtop, Institute of Physics, Polish Academy of Sciences, Aleja Lotników 32/46, 02668 Warsaw, Poland.
4Chair of Theoretical Chemistry, Technische Universität Dresden, Bergstrasse 66, 01069 Dresden, Germany.
5Departamento de Física, Universidad de Antofagasta, Av. Angamos 601, Casilla 170, Antofagasta, Chile.
The interplay of strong electronic correlations, sizable octahedral distortions, and pronounced spin–orbit coupling (SOC) makes perovskite oxides promising candidates for realizing altermagnetic phases. We study altermagnetic phases in Ca3Ru2O7, a non-centrosymmetric layered perovskite whose ground state is a Kramers-degenerate antiferromagnet. We show that an alternative Néel-type spin arrangement hosts a P-2 -wave altermagnetic state with orbital selectivity similar to Ca2RuO4. Including SOC generates a symmetry-allowed -wave component and yields a hybrid /-wave altermagnetic order. We further demonstrate that biaxial strain tunes both magnetic stability and band splitting: compressive strain beyond 2% favors the altermagnetic phase over the antiferromagnetic ground state, while tensile strain increases altermagnetic splittings by up to 9%. To quantify these trends, we define an altermagnetic figure of merit and trace its strain dependence to changes in electronic localization and octahedral geometry in this polar metal.
Introduction
11footnotetext: * Corresponding author: andrea.leon@uchile.cl22footnotetext: † jhon.gonzalez@uantof.clAltermagnets are a significant new magnetic material class, advancing fundamental research and spintronic applications [1, 2]. At the heart of altermagnetism (AM) lies crystal symmetry, which dictates the unique spin-splitting behavior. This inherent link makes strain a powerful tool for manipulating and controlling AM order in correlated materials, a frontier of active research [3, 4]. The potential of strain engineering has been recently demonstrated by the induced phase transition from an antiferromagnetic to an altermagnetic state in ReO2 [4] and the discovery of an elasto-Hall conductivity in d-wave altermagnets [5]. The physics is further enriched by spin-orbit coupling (SOC), which, by breaking time-reversal symmetry, can induce phenomena such as weak ferromagnetism through spin canting or other mechanisms dependent on crystal symmetry [6, 7, 8, 9, 10, 11, 12]. Collectively, these findings motivate the exploration of strain-induced AM order in new classes of correlated materials. In this context, Ruddlesden-Popper perovskites, renowned for hosting diverse quantum phases intrinsically coupled to lattice distortions, offer a compelling platform to investigate the interplay between strain, symmetry, and magnetism [13, 14, 15, 16, 17].
Ruddlesden-Popper compounds of the form (Ca, Sr)n+1RunO3n+1 exhibit strong electronic correlations and display phases that depend critically on symmetry and octahedral distortions. Rotations, Jahn-teller distortions, and polar displacements break the operations that interchange opposite spin sublattices and enable the emergence of altermagnetic states [16]. Over the past decade, researchers have observed Mott insulating behavior [18], metamagnetism [19], and unconventional superconductivity [20] in these systems. Revisiting their magnetic phases within the altermagnetic framework offers new insights into transitions among ferromagnetism, antiferromagnetism, and altermagnetism by linking symmetry, distortion, and magnetic order. Among the series, Ca3Ru2O7 undergoes antiferromagnetic transitions as a function of temperature [21, 22, 23] and exhibits pronounced tunability under magnetic field [24], pressure [25] and strain [26]. These observations raise the question of whether Ca3Ru2O7 (hereafter CRO) hosts an altermagnetic phase and what conditions might induce a transition from its native antiferromagnetic order.
We apply this framework to Ca3Ru2O7, a non-centrosymmetric compound with strong spin-orbit coupling (SOC) that enables relativistic spin splitting and spin-momentum locking. SOC governs the magnetic anisotropy in this system and drives spin reorientation transitions, as previously reported [27]. Investigating altermagnetism in Ca3Ru2O7 offers a unique opportunity to reveal spin textures that emerge from the interplay of non-relativistic and relativistic effects. Most known altermagnets preserve inversion symmetry; finding altermagnetic order in a non-centrosymmetric metal would open new directions in the study of symmetry-protected spin transport phenomena [28, 29].
In this work, we identify magnetic configurations in Ca3Ru2O7 that host altermagnetic order and analyze their symmetry and electronic structure under non-relativistic and relativistic spin splitting. In the absence of SOC, the system exhibits a -wave altermagnetic state, while SOC introduces additional asymmetries and generates a hybrid - and -wave character. We demonstrate that biaxial strain tunes the altermagnetic response by enhancing or suppressing spin splitting depending on the strain direction and link this manipulation to changes in octahedral distortions and orbital localization. To quantify these effects, we define the altermagnetic quantity (AMQ) as a figure of merit that captures changes in band splitting and enables direct comparison across altermagnetic systems across different conditions.
Results
Exploring AM states
@@ Results
Ca3Ru2O7 crystallizes in the non-centrosymmetric space group Bb21m [22], with associated point group mm2, which includes a twofold rotation axis along and a vertical mirror plane. The full space group also contains a 21 screw axis along , which plays a role in the nonsymmorphic symmetries relevant to the altermagnetic configuration. [30]. The magnetic moment arises predominantly from the Ru 4 electrons, in which, at low temperatures, the Ru atoms order ferromagnetically within the bilayers while aligning antiferromagnetically between them, forming an A-type antiferromagnetic structure where the spins are oriented along the -axis (see Fig. 1). Given this A-type antiferromagnetic order, the primitive cell of CRO contains spin-up sites, with a translational symmetry operation that maps spin-up into spin-down sites (along the -axis). Consequently, the ground state is not altermagnetic (see Fig. 1) [31].
@@
We analyzed CRO under different magnetic configurations to explore possible AM states and their stability. Specifically, we examined magnetic orderings commonly observed in AB3 perovskites [33], labeled as configurations A, B, C, and D (see Fig. 2). Table 1 shows the energy differences relative to the ground-state configuration A, as well as some relevant structural, electronic, and magnetic properties for our study. Our results indicate that AM states appear in configurations B and C, where configuration B has an antiferromagnetic spin arrangement along the - and -axes, while configuration C shows this along the -axis. Although these magnetic states have not been experimentally observed in CRO, similar phases have been reported in doped Ca3Ru2O7 systems, such as Ti-doped Ca3(Ru1-xTix)2O7 (configuration B) [34], and in the sister compound Sr3Ru2O7 (configuration D) [35].
@@
Notably, variations in magnetic spin arrangements lead to significant structural and electronic changes (further structural details are given in supplementary information file (SI) table S1). Specifically, FM and AFM in-plane arrangements result in metallic (A, C) and a narrow insular state (B, D), respectively. Next, we will focus on the AM phases of CRO, primarily in Configuration B.
@@
In addition to the electronic differences, configuration B exhibits a lower magnetic moment than configuration A (details on Table S1).
@@
In configuration B, the magnetic moments form a Néel-type collinear ordering with alternating spin orientations. This promotes spin-compensated orbital overlap and wave function delocalization, which are characteristic of bonding-like states. As a result, configuration B is more sensitive to the Hubbard- term due to the enhanced electronic proximity.
@@
As a result, a band gap opens at relatively low U values, specifically . In contrast, configuration A remains metallic under the same conditions. The Supplemental Material (Section SII) provides further analysis.
@@
We now examine the electronic structure of both configurations. Figure 3 shows the total and orbital-resolved density of states. While configuration A is metallic, configuration B exhibits a narrow gap at the Fermi level. In an ideal octahedral environment, the Ru , , and orbitals are degenerate. Orthorhombic distortions lift this degeneracy slightly, but it is the Néel-type magnetic order in the -plane (conf. B) that drives orbital differentiation by enhancing -type hybridization between Ru- and O- orbitals. This leads to a strong bonding–antibonding splitting, pushing the bonding states to lower energy and making them nearly filled. The resulting spectral redistribution lowers the chemical potential relative to the less-affected, half-filled and orbitals. The gap thus reflects an orbital-selective mechanism driven by the interplay between magnetism and crystal-field effects, as proposed for Ca2RuO4 [40]
@@
Figure 4 shows the non-relativistic band structure for the A and B magnetic configurations. The A configuration (ground state) exhibits metallic behavior, consistent with previously reported results that neglect SOC effects [27, 36]. In contrast, configuration B displays a narrow gap and a -dependent non-relativistic spin splitting, a hallmark of altermagnetic states (see Fig. 4(b)). The shaded regions on the band structure indicate regions in -space where the altermagnetic spin splitting is maximal. Figure 5 shows the detailed Brillouin zone (BZ) corresponding to these paths.
@@
The presence of the non-relativistic spin-splitting is independent of the value of ky. The observed spin splitting exhibits what is known as a d-wave symmetry in momentum space. This terminology refers to how the splitting changes sign across the Brillouin zone, in analogy to the lobes of a d-orbital. Specifically, the splitting changes sign upon inversion of the kx or kz coordinate, but remains unchanged when inverting ky. This independence from the ky component allows for a more precise classification as a dxz-wave character. According to the formal symmetry classification of altermagnets [31], this behavior corresponds to a P-2 dxz-wave state. These paths, which extend beyond the xy plane, have not been explored in previous studies, where ARPES experiments primarily focused on the xy plane [27, 41]. The following section discusses the preference for a specific direction in the AM character.
@@
The bands located between eV exhibit significant non-relativistic spin splitting, corresponding to the - orbitals. In contrast, the bands between eV, dominated by the orbital character, show negligible spin splitting (see, Fig. S3). Therefore, configuration B exhibits orbital-selective altermagnetism with the altermagnetism present in the - bands but not in the bands, this is similar to that observed in Ca2RuO4 [40]. When the or coordinate is inverted in reciprocal space, the non-relativistic spin splitting changes sign, while inversion of leaves it unchanged. Notably, this altermagnetic order matches that found in Ca2RuO4, despite its crystallization in a different space group [40]. The orbital-selective magnetism depends on the magnetic space group and orbital character. In conf.B, the nearly filled dxy orbitals do not contribute to magnetism, leading to orbital-selective AM. Consequently, conf. C displays different magnetic states in which all the t2g orbitals contribute to the AM character; therefore, it does not exhibit orbital selective AM, see more details in Fig. S3.
The lack of inversion symmetry in Ca3Ru2O7 enables sizable SOC effects, which play a central role in its magnetic properties [27]. In configuration B, as in configuration A, the spins align primarily along the -axis, with a slight canting toward -axis (). SOC enriches the spin textures and activates symmetry-breaking mechanisms that lift band degeneracies, thereby enabling access to AM states that remain hidden in the collinear case.
Our spin-projected band structure analysis reveals how SOC qualitatively transforms the nature of altermagnetism in this system. The non-relativistic ground state exhibits a clear -wave character, dominated by -like states along the high-symmetry paths shown in Fig. 4. Upon including SOC, a complex landscape emerges. The original -wave features remain most evident in the projection (Figs. 6(a)-(b)), consistent with the magnetic anisotropy along the -axis. In stark contrast, the projection unveils a completely new altermagnetic signature between -0.5 and -0.75 eV, emerging from bands that were purely antiferromagnetic in the collinear case (Figs. 6(a),(c)). The influence of SOC is particularly evident along the direction. Here, SOC lifts the band degeneracy in the channel, creating a spin splitting where the non-relativistic bands were degenerate (Figs. 6(e)-(f)) (similar results are found for along ). This purely relativistic splitting exhibits a characteristic antisymmetric behavior for momentum inversion (), which is the defining hallmark of a -wave altermagnetic component.
Therefore, these findings demonstrate the emergence of a novel hybrid -wave altermagnetic order. The AM+SOC phase in CRO is best described as a state where the non-relativistic -wave component coexists with a purely SOC-driven -wave character. This hybridization arises from the interplay between the magnetic structure and the crystal’s nonsymmorphic symmetries.
While these results highlight a recently emerging regime of altermagnetism in the presence of SOC [42, 43], a formal classification of this hybrid order, for instance, through advanced group-theoretical analysis, remains a substantial theoretical challenge. we describe the emergent AM+SOC features through their direct, symmetry-resolved signatures in the spin-projected electronic structure. Future studies will address a more systematic classification of AM+SOC spin textures.
Symmetry analysis
@@
In summary, the operation z becomes necessary (Fig. 7(f)). These combined nonsymmorphic operations result in the -dependent splitting states [31]. Due to the symmetry operations that connect the spins, which are determined by rotations and translations along the and directions while leaving the direction unchanged, the AM character is found along these axes, as shown in Fig. 4. In the case of the C configuration, the symmetry operation ( +, + , - + ) becomes necessary to achieve AM order. This symmetry operation is specific to the orthorhombic space group Bb21m. Since the crystallographic axes are not symmetry-equivalent, this operation does not remain valid under the replacement
Tuning AFM to AM phases
@@
Next, we investigate the effects of compressive (-) and tensile (+) strain on the AM features of CRO. Fig. 9 shows the band structure of configuration B under strains of = -2%, 0, and 2% (biaxial). It can be observed that compressive strain suppresses the AM features among the energy range [-2:-1] eV (compare Fig. 9 (a) with Fig. 9(b)). Moreover, the tensile strain enhances the splitting and shifts it to higher energies.
@@
Note that the integration to obtain the AMQ can be performed either over the entire Brillouin zone or along a specific -path. We verified that both methods yield similar solutions; the only difference is a constant multiplicative factor, which cancels out when computing percentage changes.
We sum up all the occupied states rather than comparing only a few selected bands. This approach proves more reliable because the most strongly split bands can vary across different systems and even within the same material under strain, depending on details such as band character and crystal symmetry. Consequently, integrating over the entire set of occupied bands offers a more robust assessment of the altermagnetic spin-splitting.
@@
in our code, Simpson’s rule is used for numerical integration. When points are expressed in inverse angstroms (), the resulting AMQ is in units of . Although the AMQ is not a formal symmetry-derived order parameter like magnetic multipoles [44], it provides a symmetry-consistent and practical metric to quantify momentum-resolved spin splitting in systems with zero net magnetization. While this approach is robust, it may fail near band crossings or overlapping bands. A more rigorous treatment, such as disentanglement via Wannier-based orbital tracking, could refine the analysis but lies beyond the scope of this study. Nonetheless, the observed vanishing of spin splitting at high-symmetry points and the overall internal consistency strongly support the identification of altermagnetic behavior.
@@
Fig. S6 shows the evolution of the spin splitting across the full -path and occupied bands under different strain conditions (see Brillouin zone in Fig. 4). The magnitude of band-dependent altermagnetic splitting varies with strain, increasing notably under tensile strain. The AQM response varies by approximately under tensile strain (2%) and under compressive strain (-2%), as shown in Fig. 8(c). All computational details and equations are provided in the Supplementary Information (Fig. S7).
Discussion
The non-relativistic spin-splitting in CRO arises from the interplay between hopping parameters and the opposite on-site spin polarizations in the two magnetic sublattices. Strain introduces competing effects that influence these parameters: compression along one axis promotes electronic delocalization, weakening magnetism and reducing the AM band splitting, whereas tensile strain favors localization, enhancing magnetic exchange and increasing the AM band splitting.
From the energy-strain relationship shown in Fig. 8(a), we observe that configuration B (the AM phase) becomes energetically more favorable than configuration A (the AFM phase) under biaxial compressive strain (). Compressive strain decreases the in-plane Ru-O bond lengths, forcing an expansion along the -axis. This expansion modifies the balance of structural distortions and stabilizes configuration B over configuration A. At the electronic level, the in-plane compression broadens the bands, weakening the magnetic exchange interactions and consequently reducing the AM band splitting. Conversely, under tensile strain, the lattice is stretched in the -plane while slightly contracting along , effectively narrowing the bandwidth and reinforcing the altermagnetic character (Fig. S5). This effect is reflected in the rising AMQ values in Fig. 8(c) RS line (relaxed structures), demonstrating how strain can selectively enhance or suppress altermagnetic character.